The measure of an angle is 46 degrees less than the measure of its supplement. find the measure of the angle and the measure of the supplement.
Answer by checkley77(12844)

X+(X-46)=180
2X=180+46
2X=226
X=226/2
X=113 DEGREES FOR THE LARGER ANGLE.
180-113=67 DEGREES FOR THE LESSER ANGLE.
PROOF:
113+67=180
180=180